OpenStudy (anonymous):

hat is the solution to the equation 8-2b = - 2/3(12b + 6) ? b =

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

U can distribute 2/3 to 12b and 6 first

OpenStudy (godlovesme):

@AidanWithanAnota can first distribute 2/3 to the terms inside the parenthesis?

OpenStudy (anonymous):

I think so @Godlovesme

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

\[8-2b=-\frac{ 2 }{ 3 }(12b+6)\]

OpenStudy (dtan5457):

correct?

OpenStudy (godlovesme):

\[8-2b=8b+4\] this should be what u get after distributing 2/3 @AidanWithanAnota

OpenStudy (anonymous):

6b and 12b

OpenStudy (godlovesme):

how did u get that?

OpenStudy (dtan5457):

If the equation i typed is correct. then distribute to get: 8b-2b=-8b+4

OpenStudy (dtan5457):

Sorry *8-2b=-8b+4

OpenStudy (dtan5457):

then you just isolate B like i told you yesterday and solve for it

OpenStudy (godlovesme):

oh i didn't c the -ve sign oppps